# Optimizing Large Language Models for ESG Activity Detection in Financial Texts

### v1

@@ IV-C Experimental Setup
In order to study the advantages of fine-tuning LLMs for ESG activity classification, we evaluated the models in two different modes: zero-shot and fine-tuning. The fine-tuning was performed on the two alternative training sets within ESG-activities: the manually curated dataset of 212 instances and a larger dataset of 1,272 instances, which combines original and synthetic data. In the following sections, we provide a detailed discussion of both approaches.
IV-C1 Zero-shot
@@ IV-C2 Fine-tuning
In this study, we fine-tuned eight LLMs for the proposed task using the training and validation datasets from the ESG-Action benchmark. Specifically, each model was fine-tuned twice: first using the manually curated training and validation datasets, and then using augmented datasets that included synthetic data. This approach enables a comparative analysis of the two model versions, allowing us to evaluate the impact of synthetic data on models of different sizes.
@@
The optimization process was handled using the AdamW optimizer [29], a variant of the Adam optimizer that incorporates decoupled weight decay regularization. To optimize the fine-tuning process, we employed LoRA techniques from the PEFT library, as previously mentioned. This approach reduces the number of trainable parameters while preserving model performance. For instance, when fine-tuning Llama 3B, which has over 3.3 billion parameters, LoRA required modifying only 0.08% of the total parameter space. Similarly, for Llama 7B, only 0.12% of its parameters were updated.
The only exception to this procedure was GPT-4o Mini, which we fine-tuned using Azure OpenAI [^14] Azure OpenAI - https://azure.microsoft.com/en-us/products/ai-services/openai-service, a cloud-based service provided by Microsoft that enables access to OpenAI’s language models. We formatted the training dataset in the required JSON structure, uploaded it to Azure, and executed the fine-tuning process with the default hyperparameters.
V Evaluation
In this section, we present the results of our experiments on the ESG-Activities benchmark. Specifically, we evaluated the eight LLMs described in Section IV using three alternative configurations: 1) zero-shot learning, 2) fine-tuning on the original data from ESG-Activities, and 3) fine-tuning on a combination of the original and synthetic data. All models were tested on the test set of ESG-Activities.
As the task is framed as a binary classification problem, we used precision, recall, and F1 score as evaluation metrics.
V-A Zero-Shot Setting
Table I reports the performance of the models in a zero-shot setting. GPT4o_mini yields the best result (72.7%), followed by GEMMA_7B (70.5%), and Llama3_8B (70.1%). GEMMA_2B and Llama_3B performed fairly well, exceeding 68% F1. Conversely, RecurrentGEMMA_2B and MISTRAL showed relatively poor performance in this task. This suggests that the recurrent architecture of RecurrentGEMMA_2B may not offer significant benefits in this specific task. Overall, there is considerable room for improvement.
@@ V-A Zero-Shot Setting
TABLE I: Performance of the models in a zero-shot setting.
V-B Fine-tuning
Table II presents the results of the models fine-tuned on the manually curated dataset. The results demonstrate a significant improvement across all models, confirming the efficacy of the ESG-Action benchmark in optimizing performance for this task.
The top-performing model is again GPT4o_mini, which attains an F1 score of 80.5%, followed by Llama_3B (78.9%) and GEMMA_2B (78.1%). It is interesting that the smaller model achieved the best performance after fine-tuning. This is likely due to the fact that the manually curated training set is relatively small. As a result, models tend to overfit quickly, and this effect is even more pronounced in larger models, which overfit earlier and thus experience a drop in performance.
@@ V-B Fine-tuning
TABLE II: Performance of the fine-tuned models (original data).
Table III presents the performance of models fine-tuned on the second training set of ESG-activities, which includes a combination of original and synthetic data. This experiment aimed to determine whether incorporating synthetic data during fine-tuning provides a performance advantage. The results clearly indicate that it does, with most models now achieving F1 scores exceeding 80%. The best overall result is achieved by Llama2_7B, with an F1 score of 84.9%, followed by Llama_3B at 84.4% and GEMMA_7B at 82.1%. GPT4o_mini, which was fine-tuned using the OpenAI service, did not perform as well in this setting.
@@
TABLE III: Performance of the fine-tuned models (original and synthetic data).
In conclusion, our experiments demonstrate that LLMs can achieve high performance in classifying text based on ESG activities when fine-tuned on high-quality datasets. These findings also provide further evidence supporting the effectiveness of synthetic data in specific domains. Notably, the most comprehensive version of the ESG-Activities benchmark training set, which combines both manually annotated and synthetic data, further improves classification performance. For example, Llama2_7B, the best-performing model, achieved a remarkable 14.8% increase in F1 score, rising from 70.1% to 84.9% with the inclusion of synthetic data.
V-C Training and Inference Time
Training times varied significantly across models, reflecting differences in architecture, size, and the complexity of the fine-tuning processes. Analyzing these variations offers valuable insights into the computational demands of each model.
Table IV provides the training and inference times for the various models. It was not possible to report the fine-tuning time for GPT4o_mini, as this process was managed by the OpenAI service.
@@ V-C Training and Inference Time
(s) Original (s) Original+Synthetic (s)
@@
TABLE IV: Training and inference time for the eight models.
@@ V-D Cost Analysis
TABLE V: Fine-tuning costs
@@
Table V reports the costs associated with fine-tuning the eight models on both the original dataset and the full version, which includes synthetic data. All models, except GPT4o_mini, were trained on an NVIDIA A100 GPU. The hourly rental rate for an A100 machine was approximately \$9 at the time of this study. Differently, the fine-tuning of GPT4o_mini was conducted using OpenAI’s standard API offering, which, at the time of writing, allowed fine-tuning at a cost of \$3.00 per 1 million tokens. For these reasons, although GPT4o_mini is the largest model in terms of the number of parameters, it is the cheapest to train using OpenAI’s services. However, this training cost must be combined with the fees required to deploy and maintain the model on OpenAI’s infrastructure, which is currently \$1.62 an hour. Among the open models, the most cost-effective to fine-tune with our configuration are Gemma_2B and Gemma_7B.
VI Related work
This section examines key advancements in the development and application of LLMs in the ESG domain. LLMs have demonstrated remarkable capabilities in processing textual data at scale, enabling the accurate extraction and synthesis of information from diverse sources, including research papers [30], patents [31], medical records [32], social media posts [33], news articles [34], technical documentation [35], legal texts [36], and financial documents [37, 38, 39].
Recent LLM families, including GPT [9], Llama [5], Mistral [8], and Gemma [6], have demonstrated outstanding capabilities in analyzing and classifying ESG and financial data. Their proficiency in processing vast amounts of unstructured information positions them as powerful tools for enhancing ESG reporting, regulatory compliance, and data-driven decision-making. However, the standard versions of these models, when used with simple prompting, do not always achieve optimal performance. Consequently, many researchers focus on developing domain-specific adaptations by fine-tuning open models on ESG and financial data. For instance, Yang et al. [40] introduced FinGPT, an open-source LLM designed for the financial sector. Initially trained on a general-purpose corpus, FinGPT was then fine-tuned using domain-specific datasets, including financial news, earnings reports, and real-time market data. This process enables the model to capture the nuanced complexities of financial language, improving its performance in tasks such as market forecasting, sentiment analysis, and news classification. As a result, FinGPT can be applied to various use cases, including robo-advising, algorithmic trading, and low-code development.
FinBERT [41] is a domain-specific adaptation of BERT [42] designed for financial text analysis. It is pre-trained on a vast corpus of financial documents, including regulatory filings, corporate announcements, and financial articles, before undergoing fine-tuning for specific tasks. This specialized training enables FinBERT to develop a deep understanding of financial terminology and contextual nuances. Notably, it has demonstrated exceptional performance in sentiment classification of financial analyst reports. Similarly, ESG-BERT [43] is a specialized variant of BERT that has been pre-trained on a large corpus of sustainability-related texts. This model has demonstrated strong performance in various natural language processing tasks within the ESG domain, particularly in text classification and sentiment analysis [44]. We initially explored adapting these models to the task analyzed in this paper; however, this approach proved unfeasible. These models are trained for specific classification tasks with predefined labels, making it difficult to adapt them for matching a text with an arbitrary activity description.
Several approaches have been proposed for ESG sentiment classification, which involves categorizing text as either positive/opportunity or negative/risk from an ESG perspective. Notably, the ML-ESG-2 shared task [45], co-located with the FinNLP workshop at IJCNLP-AACL 2023, has played a key role in advancing methodologies in this field. To tackle this challenge, researchers have leveraged both encoder-only and decoder-only transformer models, including fine-tuned versions of FinBERT [46] and LLMs such as Llama 2 and Dolly  [47]. Similar to previous approaches, we leverage and evaluate a variety of fine-tuned language models. However, the task addressed in this paper is more fine-grained than merely detecting a positive or negative sentiment.
To tackle the challenges of long contexts and data recency in LLMs, researchers have started incorporating the RAG paradigm [11], which improves response accuracy by dynamically retrieving relevant information. For instance, ESGReveal [47] is a recent method for extracting and analyzing ESG data from corporate reports that utilizes an LLM augmented with RAG. The system includes an ESG metadata module for precise querying, a preprocessing module for database construction, and an LLM agent for extracting information. Like our approach, ESGReveal leverages an RAG architecture. However, its primary focus is on extracting quantitative data from documents, such as Scope 1 and Scope 2 emissions. In contrast, our work is centred on identifying portions of documents that align with the activities described in the ESG taxonomy. Furthermore, ESGReveal utilizes zero-shot learning LLMs with basic prompting, whereas we fine-tune LLMs to optimize their performance for our specific task.
Efforts have also been made to develop a question-answering system in the ESG domain using a pipeline based on Knowledge Graph-based Retrieval Augmented Generation (KG-RAG) [48]. Unlike traditional RAG, which retrieves information from textual sources, KG-RAG leverages knowledge graphs [49], which are structured representations of information that capture relationships between entities in a graph format. Similarly, Angioni et al. [50] employ a combination of linguistic pattern analysis and transformer models to extract a knowledge graph from a corpus of news articles, aiming to capture key trends related to ESG aspects. In this work, we focus on the more specific task of classifying text based on ESG activities. Consequently, we rely exclusively on action descriptions from the ESG taxonomy rather than employing a more complex knowledge representation.
VII Concusion
@@
Our results demonstrate that fine-tuned models significantly outperform zero-shot approaches, confirming that domain adaptation plays a crucial role in enhancing ESG text classification. Moreover, the inclusion of synthetic data proved to be a valuable strategy, yielding substantial performance gains across multiple model architectures. This finding suggests that synthetic data augmentation can serve as an effective method to overcome data scarcity issues, particularly in specialized domains such as ESG analysis.
Another key takeaway from our experiments is that relatively small, open-source models can achieve competitive performance levels when fine-tuned with high-quality, domain-specific data. This insight has significant implications for organizations looking to deploy cost-effective AI solutions for ESG analysis or those that need to run models locally due to privacy and data security concerns, which are particularly common in the financial domain.
@@
Acknowledgment
This work is supported by the Italian Ministry of University and Research (MUR) within the PRIN2022—ISALDI: Interpretable Stock Analysis Leveraging Deep multImodal models (CUP: E53D23008150006).

### v2

@@ 4.3. Experimental Setup
In order to study the advantages of fine-tuning LLMs for ESG activity classification, we evaluated the models in two different modes: zero-shot and fine-tuning. The fine-tuning was performed on the two alternative training sets within ESG-activities: the manually curated dataset of 212 instances and a larger dataset of 1,272 instances, which combines original and synthetic data. In both cases, the models were evaluated on the ESG-activities test set using standard metrics for binary classification: precision, recall, and F1-score. We used the weighted-averaged version of these metrics, given the unbalanced nature of the dataset, which contains more instances of class 0 than class 1. In the following sections, we provide a detailed discussion of both approaches.
4.3.1. Zero-shot
@@ 4.3.2. Fine-tuning
In this study, we fine-tuned nine models for the proposed task using the training and validation datasets from the ESG-Activities benchmark. Specifically, each model was fine-tuned twice: first using the manually curated training and validation datasets, and then using augmented datasets that included synthetic data. This approach enables a comparative analysis of the two model versions, allowing us to evaluate the impact of synthetic data on models of different sizes.
@@
The optimization process was handled using the AdamW optimizer (guan2023weightpredictionboostsconvergence), a variant of the Adam optimizer that incorporates decoupled weight decay regularization. To optimize the fine-tuning process, we employed LoRA techniques from the PEFT library, as previously mentioned. This approach reduces the number of trainable parameters while preserving model performance. For instance, when fine-tuning Llama 3B, which has over 3.3 billion parameters, LoRA required modifying only 0.08% of the total parameter space. Similarly, for Llama 7B, only 0.12% of its parameters were updated.
To enhance transparency and reproducibility, we summarize the main experimental settings and hyperparameters used during the fine-tuning process in Table 1. These configurations were applied consistently across all model variants unless explicitly stated otherwise. The settings were chosen based on preliminary experiments and established best practices for low-resource fine-tuning scenarios.
Parameter Value
Gradient accumulation steps 5
Learning rate 3e-4
Validation split ratio 0.15
Cross-validation folds 10
Evaluation steps Every 10% of total steps
Optimizer AdamW
Adapter type LoRA
LoRA rank ($r$) 8
LoRA alpha 32
LoRA dropout 0.05
LoRA target modules ["query_key_value"]
Table 1. Summary of experimental and hyper-parameter settings for fine-tuning.
Two models required deviations from the standard fine-tuning procedure described above. GPT‑4o Mini was fine‑tuned using Azure OpenAI [^14] Azure OpenAI — https://azure.microsoft.com/en-us/products/ai-services/openai-service, a cloud‑based service provided by Microsoft that offers access to OpenAI’s language models. We formatted the training dataset in the required JSON structure, uploaded it to Azure, and executed the fine-tuning process with the default hyperparameters. ESG‑BERT also required an additional step. This model was originally designed as a 26‑class classifier specifically for ESG text classification tasks. Since no publicly available version of the model without the task‑specific 26‑class output layer was available, we adapted the existing checkpoint by removing its final classification layer and replacing it with a binary output layer, making it suitable for the binary classification task defined in this study.
5. Evaluation
In this section, we present the results of our experiments on the ESG-Activities benchmark. Specifically, we evaluated the nine models described in Section 4 using three alternative configurations: 1) zero-shot learning, 2) fine-tuning on the original data from ESG-Activities, and 3) fine-tuning on a combination of the original and synthetic data. As previously mentioned, we evaluated the performance of the models using standard metrics for binary classification: precision, recall, and F1-score.
5.1. Zero-Shot Setting
Table 2 reports the performance of the models in a zero-shot setting. GPT4o_mini yields the best result (72.7%), followed by Gemma_7B (70.5%), and Llama3_8B (70.1%). Gemma_2B and Llama_3B performed fairly well, exceeding 68% F1. Conversely, ESG-BERT, RecurrentGemma_2B and Mistral_7B showed relatively poor performance in this task. Overall, there is considerable room for improvement.
@@ 5.1. Zero-Shot Setting
ESG-BERT 0.6994 0.5094 0.5455
Table 2. Performance of the models in a zero-shot setting.
5.2. Fine-tuning
Table 3 presents the results of the models fine-tuned on the manually curated dataset. The results demonstrate a significant improvement across all models, confirming the efficacy of the ESG-Activities benchmark in optimizing performance for this task.
The top-performing model is again GPT4o_mini, which attains an F1 score of 80.5%, followed by Llama_3B (78.9%) and Gemma_2B (78.1%). It is interesting that some of the smaller models achieved the best performance after fine-tuning. This is likely due to the fact that the manually curated training set is relatively small. As a result, models tend to overfit quickly, and this effect is even more pronounced in larger models, which overfit earlier and thus experience a drop in performance. Notably, ESG-BERT achieves a high precision (82.7%) but suffers from a low recall, which leads to a lower overall F1 score. Several LLMs achieve better performance when fine-tuned on ESG-Activities.
@@ 5.2. Fine-tuning
ESG-BERT 0.8272 0.6415 0.6689
Table 3. Performance of the fine-tuned models (original data).
Table 4 presents the performance of models fine-tuned on the second training set of ESG-activities, which includes a combination of original and synthetic data. This experiment aimed to determine whether incorporating synthetic data during fine-tuning provides a performance advantage. The results clearly indicate that it does, with most models now achieving F1 scores exceeding 80%. The best overall result is achieved by Llama2_7B, with an F1 score of 84.9%, followed by Llama_3B at 84.4% and Gemma_7B at 82.1%. GPT4o_mini, which was fine-tuned using the OpenAI service, did not perform as well in this setting. As before, ESG-BERT underperformed.
@@
ESG-BERT 0.8360 0.6792 0.7050
Table 4. Performance of the fine-tuned models (original and synthetic data).
In conclusion, our experiments show that LLMs, particularly open lightweight models that can run in low-resource settings, achieve high performance in classifying text related to ESG activities when fine-tuned on high-quality datasets. In contrast, more traditional encoder-based approaches such as ESG-BERT are less effective. These findings also provide further evidence supporting the effectiveness of synthetic data in specific domains. Notably, the most comprehensive version of the ESG-Activities benchmark training set, which combines both manually annotated and synthetic data, further improves classification performance. For example, Llama2_7B, the best-performing model, achieved a remarkable 14.8% increase in F1 score, rising from 70.1% to 84.9% with the inclusion of synthetic data.
5.3. Training and Inference Time
Training times varied significantly across models, reflecting differences in architecture, size, and the complexity of the fine-tuning processes. Table 5 presents the training and inference times of the different models. We do not report the fine-tuning time for GPT4o_mini because it was handled by the OpenAI service.
@@ 5.3. Training and Inference Time
(s) Original (s) Original+Synthetic
@@
Table 5. Training and inference time for the nine models.
@@ 5.4. Cost Analysis
Table 6. Fine-tuning costs
@@
Table 6 reports the costs associated with fine-tuning the nine models on both the original dataset and the full version, which includes synthetic data. All models, except GPT4o_mini, were trained on an NVIDIA A100 GPU. The hourly rental rate for an A100 machine was approximately \$9 at the time of this study. Differently, the fine-tuning of GPT4o_mini was conducted using OpenAI’s standard API offering, which, at the time of writing, allowed fine-tuning at a cost of \$3.00 per 1 million tokens. For these reasons, although GPT4o_mini is the largest model in terms of the number of parameters, it is the cheapest to train using OpenAI’s services. However, this training cost must be combined with the fees required to deploy and maintain the model on OpenAI’s infrastructure, which is currently \$1.62 an hour. The open LLMs yielded similar costs, with the exception of Mistral_7B, which was clearly more expensive. ESG-BERT was the most cost-effective due to its small size; however, its performance may not be sufficient to justify the savings.
6. Related work
This section examines key advancements in the development and application of LLMs in the ESG domain. LLMs have demonstrated remarkable capabilities in processing textual data at scale, enabling the accurate extraction and synthesis of information from diverse sources, including research papers (bolanos2024artificial; cadeddu2024optimizing), patents (kosonocky2024mining), social media posts (yang2024mentallama), medical records (omiye2024large), news articles (motta2024capturing), legal texts (savelka2023unreasonable), knowledge graphs (tsaneva2025knowledge), structured representation of scientific knowledge (meloni2025exploring), and financial documents (wu2023bloomberggpt). Recent LLM families, including GPT (openai2023gpt4o), Llama (llama), Mistral (mistral), and Gemma (gemma), have demonstrated outstanding capabilities in analyzing and classifying ESG and financial data. Their proficiency in processing vast amounts of unstructured information positions them as powerful tools for enhancing ESG reporting, regulatory compliance, and data-driven decision-making. However, the standard versions of these models, when used with simple prompting, do not always achieve optimal performance. Consequently, many researchers focus on developing domain-specific adaptations by fine-tuning open models on ESG and financial data. For instance, Yang et al. (yang2023fingpt) introduced FinGPT, an open-source LLM designed for the financial sector. Initially trained on a general-purpose corpus, FinGPT was then fine-tuned using domain-specific datasets, including financial news, earnings reports, and real-time market data. Despite the availability of several versions of FinGPT, none proved suitable for our purposes. In particular, we tested a multitask version of the model configured for output generation, but it turned out to be entirely inadequate for our ESG activity classification task, as it failed to produce any concrete or usable outputs. This limitation is largely due to the nature of the ESG classification task, which requires precise, binary or categorical decisions aligned with specific regulatory taxonomies, rather than the open-ended, natural language responses that FinGPT is optimized to generate. Another domain-specific solution is ESG-BERT (mukherjee2020esg), a specialized variant of BERT that was discussed in Section 4.2 and included in our experiments. This model achieved subpar performance on the task addressed in this paper, mainly because of its low recall. Finally, FinBERT (huang2022finbert) is a domain-specific adaptation of BERT (kenton2019bert) designed for financial text analysis. It is pre-trained on a vast corpus of financial documents, including regulatory filings, corporate announcements, and financial articles, before undergoing fine-tuning for specific tasks. We attempted to adapt FinBERT for our binary classification task by removing its task-specific output layer and replacing it with a binary classification head, following the same approach applied to ESG-BERT. However, no publicly available checkpoint without a classification head could be found, and this adaptation resulted in very poor performance on ESG-Activities (F1 ¡ 20%). Based on these results, we decided not to include FinBERT in the final performance tables. Several approaches have been proposed for ESG sentiment classification, which involves categorizing text as either positive/opportunity or negative/risk from an ESG perspective. Notably, the ML-ESG-2 shared task (kannan2023textual), co-located with the FinNLP workshop at IJCNLP-AACL 2023, has played a key role in advancing methodologies in this field. To tackle this challenge, researchers have leveraged both encoder-only and decoder-only transformer models, including fine-tuned versions of FinBERT (mishra2023esg) and LLMs such as Llama 2 and Dolly  (zou2023esgreveal). Similar to previous approaches, we leverage and evaluate a variety of fine-tuned language models. However, the task addressed in this paper is more fine-grained than merely detecting a positive or negative sentiment. To tackle the challenges of long contexts and data recency in LLMs, researchers have started incorporating the RAG paradigm (lewis2020retrieval), which improves response accuracy by dynamically retrieving relevant information. For instance, ESGReveal (zou2023esgreveal) is a recent method for extracting and analyzing ESG data from corporate reports that utilizes an LLM augmented with RAG. The system includes an ESG metadata module for precise querying, a preprocessing module for database construction, and an LLM agent for extracting information. Like our approach, ESGReveal leverages an RAG architecture. However, its primary focus is on extracting quantitative data from documents, such as Scope 1 and Scope 2 emissions. In contrast, our work is centred on identifying portions of documents that align with the activities described in the ESG taxonomy. Furthermore, ESGReveal utilizes zero-shot learning LLMs with basic prompting, whereas we fine-tune LLMs to optimize their performance for our specific task. Efforts have also been made to develop a question-answering system in the ESG domain using a pipeline based on Knowledge Graph-based Retrieval Augmented Generation (KG-RAG) (gupta2024knowledge). Unlike traditional RAG, which retrieves information from textual sources, KG-RAG leverages knowledge graphs (peng2023knowledge), which are structured representations of information that capture relationships between entities in a graph format. Similarly, Angioni et al. (angioni2024exploring) employ a combination of linguistic pattern analysis and transformer models to extract a knowledge graph from a corpus of news articles, aiming to capture key trends related to ESG aspects. In this work, we focus on the more specific task of classifying text based on ESG activities. Consequently, we rely exclusively on action descriptions from the ESG taxonomy rather than employing a more complex knowledge representation.
7. Conclusion
@@
Our results demonstrate that fine-tuned models significantly outperform zero-shot approaches, confirming that domain adaptation plays a crucial role in enhancing ESG text classification. Moreover, the inclusion of synthetic data proved to be a valuable strategy, yielding substantial performance gains across multiple model architectures. This finding suggests that synthetic data augmentation can serve as an effective method to overcome data scarcity issues, particularly in specialized domains such as ESG analysis. Another key takeaway from our experiments is that relatively small, open-source models can achieve competitive performance levels when fine-tuned with high-quality, domain-specific data. This insight has significant implications for organizations looking to deploy cost-effective AI solutions for ESG analysis or those that need to run models locally due to privacy and data security concerns, which are particularly common in the financial domain.
@@
Acknowledgements.
This work is supported by the Italian Ministry of University and Research (MUR) within the PRIN2022-ISALDI: Interpretable Stock Analysis Leveraging Deep Multimodal Models (CUP: E53D23008150006).